OpenStudy (anonymous):

Find the derivative of f(x) = -9/x at x = 6.

OpenStudy (anonymous):

is it 3/2

OpenStudy (psymon):

Nah. Ill let mebs finish, though :P

OpenStudy (anonymous):

Write the equation as so .\[Y = -9(x)^{-1}\] than Use the first principle rule. \[\frac{dy}{dx} = -9 (-1)(x)^{-1 -1}\]

OpenStudy (anonymous):

so its 9x^-1-1

OpenStudy (anonymous):

Than your equation looks like this \[\frac{dy}{dx} = \frac{ 9 }{ x^{2} }\] than evaluate at x = 6.

OpenStudy (anonymous):

so its 1/4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i have three more for anyone else willing to help ill post them in a new question

OpenStudy (anonymous):

good idea.. let psymon try.